OpenStudy (anonymous):

-5(5n+2)=30 n=-8/5 please shoe me how i got this answer

OpenStudy (anonymous):

first you multiply -5 by what is in the bracket... -5(5n+2) = -25n -10..and that is to the left of your "=" sign So, you now have -25n-10 = 30 you want "25n" by itself on the left side, so you add 10 to each side 25n = 30 +10 25n = 40 Next, you want to solve for "n" so you divide each side by 25 n = 40/25 n= 8/5

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

your welcome!

OpenStudy (anonymous):

by any way what about showing work for 3(6z-1)-5(z+3)=5(z+1) z=23/8

OpenStudy (anonymous):

ok, sure, you use the same method as I showed you up above... First on the left of the "=", you want to multiply 3 by the bracket next to it and multiply -5 by the bracket beside it...

OpenStudy (anonymous):

ok i am so trying to understand this all

OpenStudy (anonymous):

18z - 3 - 5z -15 = 5z + 1 As you notice, on the right of the "=" sign I multuplied 5 by the bracket next to it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

once you have 18z - 3 - 5z -15 = 5z + 1 you can combine any "like" terms on each side 13z - 18 = 5z + 1 now, you want your "z's" on one side of the equl sign and your numbers on the other

OpenStudy (anonymous):

8z = 19 z = 19/8 what did think the answer was?

OpenStudy (anonymous):

2.38

OpenStudy (precal):